# Smooth orbit equivalence rigidity for dissipative geodesic flows

### v1

@@ Appendix A Solutions to the transport equation as extensions
A key ingredient that we use in this paper is Theorem 2.15: it allows us to extend any holomorphic $1$-form $\tau$ on $M$ (seen as a function on $SM$) to a fiberwise holomorphic distribution $u\in\mathcal{D}^{\prime}(SM)$ satisfying the transport equation $(F+V(\lambda))u=0$. We say that the distribution $u$ is an extension of $\tau$ in the sense that $u_{1}=\pi_{1}^{*}\tau$.
This can be seen as part of a larger theme, which is to find distributional solutions of the transport equation $(F+V(\lambda))u=0$ with some prescribed Fourier modes. The problem is closely related to the study of the surjectivity of the adjoint of the X-ray transform for thermostats, which in turn is key to understanding the injectivity of the X-ray transform operator itself.
As an example, the following extension result generalizes [Ain15, Theorem 1.6] and [AZ17, Theorem 1.7], which cover the cases of magnetic flows and Gaussian thermostats respectively.
@@
Let $(M,g,\lambda)$ be an Anosov generalized thermostat. For any $f\in\mathcal{C}^{\infty}(M)$, there exists $u\in H^{-1}(SM)$ such that $(F+V(\lambda))u=0$ and $u_{0}=\pi^{*}f$.
The argument crucially relies on the Pestov identity (see [DP07, Theorem 3.3]).
@@
Let $(M,g,\lambda)$ be a generalized thermostat. Then
@@
Recall that the thermostat curvature $\mathbb{K}$ for a generalized thermostat is defined in (1.4). In both [Ain15] and [AZ17], the proofs introduce the following property:
@@
Let $\alpha\in[0,1]$. A generalized thermostat $(M,g,\lambda)$ is $\alpha$-controlled if
@@
They then show that magnetic flows and Gaussian thermostats are $\alpha$-controlled for some $\alpha>0$ whenever they are Anosov. However, using the Pestov identity (A.1) for generalized thermostats, the same proof as in [AZ17, Theorem 3.1] goes through for the more general case.
@@
Let $(M,g,\lambda)$ be an Anosov generalized thermostat. Then, there exists $\alpha>0$ such that
@@
The next theorem is again in the spirit of extending functions with low Fourier degree: it deals with functions induced from 1-forms on $M$. The result requires a technical condition, which is that the smooth $1$-form $\theta$ being considered needs to be solenoidal (or divergence-free) in the sense that $\delta\theta=0$. Here $\delta$ is the co-differential with respect to the metric $g$ acting on $1$-forms, i.e., $\delta=-\star d\star$. If we write $\pi_{1}^{*}\theta=a_{-1}+a_{1}\in\Omega_{-1}\oplus\Omega_{1}$, then $\delta\theta=0$ if and only if $\eta_{+}a_{-1}+\eta_{-}a_{1}=0$ (see [PSU14]).
@@
Let $(M,g,\lambda)$ be an Anosov generalized thermostat. For any solenoidal smooth $1$-form $\theta$ on $M$, there exists $u\in H^{-1}(SM)$ with $u_{0}=0$ such that $(F+V(\lambda))u=0$ and $u_{-1}+u_{1}=\pi_{1}^{*}\theta$.
This is a generalization of [Ain15, Theorem 1.7] and [AZ17, Theorem 1.8], which again deal with the magnetic and Gaussian thermostat cases respectively. However, adapting the proofs requires some care. We will need the following lemma.
@@
For any $\lambda\in\mathcal{C}^{\infty}(SM)$, the $L^{\infty}$-norms of its Fourier coefficients $\{\lambda_{k}\}_{k\in\mathbb{Z}}$ are rapidly decaying in the sense that, for all $\alpha\in\mathbb{N}$, we have
@@
For any point on $SM$, let $U\subseteq SM$ be an open neighborhood admitting smooth coordinates $(x,\theta)\in\mathbb{R}^{2}\times\mathbb{S}^{1}$ such that $V\equiv\partial_{\theta}$. The Sobolev embedding theorem gives us a constant $C>0$ such that
$\|\lambda_{k}\|_{L^{\infty}(U)}\leq C\sum_{|\beta|\leq 2}\|D^{\beta}\lambda_{k}\|_{L^{2}(U)}$
for all $k\in\mathbb{Z}$. Here we use the multi-index notation $\beta=(\beta_{1},\beta_{2},\beta_{3})\in\mathbb{N}^{3}$ and define $D^{\beta}:=\partial_{x_{1}}^{\beta_{1}}\partial_{x_{2}}^{\beta_{2}}\partial_{\theta}^{\beta_{3}}$.
@@
Since the $L^{2}$-norms of the Fourier coefficients of a smooth function are rapidly decaying, we obtain the desired result.
The following lemma has the same proof as [AZ17, Theorem 1.8]. The argument relies on the Pestov identity and Theorem A.4.
@@
Let $(M,g,\lambda)$ be an Anosov generalized thermostat. Then there exists a constant $C>0$ such that
$\|u\|_{H^{1}(SM)}\leq C\|VFu\|$
for all $u\in\bigoplus_{|k|\geq 1}\Omega_{k}$
@@
$Tu:=\sum_{|k|\geq 2}u_{k}.$
We also define $Q:\mathcal{C}^{\infty}(SM)\to\bigoplus_{|k|\geq 2}\Omega_{k}$ as $Q:=TVF$.
@@
Let $(M,g,\lambda)$ be an Anosov generalized thermostat. Then there exists a constant $C>0$ such that
$\|u\|_{H^{1}(SM)}\leq C\|Qu\|$
@@
By the triangle inequality, the Cauchy-Schwarz inequality, Lemma A.6, and property (A.2), we obtain
@@
$\|\eta_{-}u_{2}\|\leq C\|Qu\|\quad\text{ and }\quad\|\eta_{+}u_{-2}\|\leq C\|Qu\|.$

### v2

@@ Appendix A Solutions to the transport equation as extensions
A key ingredient that we use in this paper is Theorem 2.15: it allows us to extend any holomorphic $1$-form $\tau$ on $M$ (seen as a function on $SM$) to a fibrewise holomorphic distribution $u\in\mathcal{D}^{\prime}(SM)$ satisfying the transport equation $(F+V\lambda)u=0$. We say that the distribution $u$ is an extension of $\tau$ in the sense that $u_{1}=\pi_{1}^{*}\tau$.
This can be seen as part of a larger theme, which is to find distributional solutions of the transport equation $(F+V\lambda)u=0$ with some prescribed Fourier modes. The problem is closely related to the study of the surjectivity of the adjoint of the X-ray transform for thermostats, which in turn is key to understanding the injectivity of the X-ray transform operator itself.
As an example, the following extension result generalizes [PSU14, Theorem 1.4], [Ain15, Theorem 1.6] and [AZ17, Theorem 1.7], which cover the cases of geodesic flows, magnetic systems and Gaussian thermostats, respectively.
@@
Let $(M,g,\lambda)$ be an Anosov thermostat. For any $f\in\mathcal{C}^{\infty}(M)$, there exists $u\in H^{-1}(SM)$ such that $(F+V\lambda)u=0$ and $u_{0}=\pi^{*}f$.
Their arguments crucially rely on the Pestov identity (see [DP07, Theorem 3.3]).
@@
Let $(M,g,\lambda)$ be a thermostat. Then, we have
@@
Recall that the thermostat curvature $\mathbb{K}$ is defined in equation (1.4). In all three papers, the proofs also introduce the following property.
@@
Let $\alpha\in[0,1]$. A thermostat $(M,g,\lambda)$ is $\alpha$-controlled if
@@
They then show that geodesic flows, magnetic systems and Gaussian thermostats are $\alpha$-controlled for some $\alpha>0$ whenever they are Anosov. However, using the Pestov identity (A.1) for thermostats, the same proof as in [AZ17, Theorem 3.1] goes through for the more general case.
@@
Let $(M,g,\lambda)$ be an Anosov thermostat. There exists $\alpha>0$ such that
@@
The next theorem is again in the spirit of extending functions with low Fourier degree: it deals with functions induced from 1-forms on $M$. The result requires a technical condition, which is that the $1$-form $\theta$ being considered needs to be solenoidal in the sense that $\delta\theta=0$. Here, $\delta$ is the co-differential with respect to the metric $g$ acting on $1$-forms, that is, $\delta=-\star d\star$. If we write $\ell_{1}\theta=a_{-1}+a_{1}\in\Omega_{-1}\oplus\Omega_{1}$, then $\delta\theta=0$ if and only if $\eta_{+}a_{-1}+\eta_{-}a_{1}=0$ (see [PSU14]).
@@
Let $(M,g,\lambda)$ be an Anosov thermostat. If $a\in\Omega_{-1}\oplus\Omega_{1}$ satisfies $\eta_{+}a_{-1}+\eta_{-}a_{1}=0$, then there exists $u\in H^{-1}(SM)$ with $u_{0}=0$ such that $(F+V\lambda)u=0$ and $u_{-1}+u_{1}=a$.
This is a generalization of [PSU14, Theorem 1.5], [Ain15, Theorem 1.7] and [AZ17, Theorem 1.8], which again deal with geodesic flows, magnetic systems and Gaussian thermostats, respectively. While the condition $u_{0}=0$ is not explicitly stated there, it follows directly from their proofs, as they construct solutions of the form $u=VTh+a$ for some $h\in L^{2}(SM)$, where the projection operator $T$, defined below in equation (A.2), kills the Fourier modes of degree $\leq 1$.
However, adapting the proofs requires some care. We will need the following lemma.
@@
For any $\lambda\in\mathcal{C}^{\infty}(SM)$, the $L^{\infty}$ norms of its Fourier modes $\{\lambda_{k}\}_{k\in\mathbb{Z}}$ are rapidly decaying in the sense that, for all $\alpha\in\mathbb{N}$, we have
@@
For any point on $SM$, let $U\subset SM$ be an open neighbourhood admitting smooth coordinates $(x,\theta)\in\mathbb{R}^{2}\times\mathbb{S}^{1}$ such that $V=\partial_{\theta}$. The Sobolev embedding theorem gives us a constant $C>0$ such that
$\|\lambda_{k}\|_{L^{\infty}(U)}\leq C\sum_{|\beta|\leq 2}\|D^{\beta}\lambda_{k}\|_{L^{2}(U)}\qquad\text{for all }k\in\mathbb{Z}.$
Here, we use the multi-index notation $\beta=(\beta_{1},\beta_{2},\beta_{3})\in\mathbb{N}^{3}$ and $D^{\beta}\coloneqq\partial_{x_{1}}^{\beta_{1}}\partial_{x_{2}}^{\beta_{2}}\partial_{\theta}^{\beta_{3}}$.
@@
Since the $L^{2}$ norms of the Fourier modes of a smooth function are rapidly decaying, we obtain the desired result. ∎
The following lemma has the same proof as [AZ17, Lemma 4.1]. The argument relies on the Pestov identity and Theorem A.4.
@@
Let $(M,g,\lambda)$ be an Anosov thermostat. Then, there exists a constant $C>0$ such that
$\|u\|_{H^{1}(SM)}\leq C\|VFu\|_{L^{2}(SM)}$
for all $u\in\bigoplus_{|k|\geq 1}\Omega_{k}$.
@@
$Tu\coloneqq\sum_{|k|\geq 2}u_{k}.$ (A.2)
We also define $Q:\mathcal{C}^{\infty}(SM)\to\bigoplus_{|k|\geq 2}\Omega_{k}$ as $Q\coloneqq TVF$.
@@
Let $(M,g,\lambda)$ be an Anosov thermostat. Then, there exists a constant $C>0$ such that
$\|u\|_{H^{1}(SM)}\leq C\|Qu\|_{L^{2}(SM)}$
@@
By the triangle inequality, the Cauchy–Schwarz inequality, Lemma A.6 and property (A.3), we obtain
@@
$\|\eta_{-}u_{2}\|\leq C\|Qu\|\qquad\text{and}\qquad\|\eta_{+}u_{-2}\|\leq C\|Qu\|.$